OpenStudy (anonymous):

6 and 1/4gsllons = ?pints

OpenStudy (anonymous):

6 and 1/4 what = ? pints ??? Is it gallons, or liters, or what ?

OpenStudy (anonymous):

6.25 gallons = x pints....1 gallon = 8 pints Set it up as a proportion... 1/8 = 6.25/x (1 gal to 8 pints = 6.25 gallons to x pints) Now cross multiply... (1)(x) = (8)(6.25) Now solve for x and that will be your answer. Can you take it from here ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

glad to help :)